OpenStudy (anonymous):

A truck is moving at a rate of 105 kilometers per hour, and the diameter of its wheels is 1 meter. Find the angular speed of the wheels in radians per minute.

OpenStudy (anonymous):

w=theta/r

OpenStudy (anonymous):

Ok, so this is confusing to me because I don't know where to start. I don't know how many revoluitions.

OpenStudy (anonymous):

find V in m/s r=1/2 V=r*omega => omega =V/r

OpenStudy (anonymous):

hmm. sorry, I don't think we use that formula.

OpenStudy (anonymous):

I mean the omega part...

OpenStudy (anonymous):

omega is angular velocity.

OpenStudy (anonymous):

Angular speed =theta/t

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

OK we use angular speed=theta/time

OpenStudy (anonymous):

k.

OpenStudy (anonymous):

do I have to convert the kilometers to miles?

OpenStudy (anonymous):

also remember once you find omega to convert from rad.s / sec to rad.s /min , since that's the units they want (for no apparent reason)

OpenStudy (anonymous):

meters / second or meters/ minute, I guess.

OpenStudy (anonymous):

well...on the omega...we haven't used that term. So I don't know how to do that.

OpenStudy (anonymous):

Here's an example of what I think I am supposed to do.

OpenStudy (anonymous):

The blades of a wind turbine are 116 feet long. The propeller rotates at 15 revolutions per min.

OpenStudy (anonymous):

Maybe that's a bad example.

OpenStudy (anonymous):

there's nothing to do. just like 'x' is distance and x/t is velocity (for unaccelerated motion) "theta" is 'angular distance' and theta/t is angular velocity

OpenStudy (anonymous):

a. Because each relvoltion generates 2pi radians, it follows that the propeller turns (15)(2pi)=30pi radians per min. The angular speeds is

OpenStudy (anonymous):

Angular speed =theta/r

OpenStudy (anonymous):

30pi radians/1 min=30 radians per min

OpenStudy (anonymous):

I don't want to sound like a typical cry baby in math..but the work give one example then the homework is similar but....it's a vague book.

OpenStudy (anonymous):

So, given that's the only example in the section...IDK.

OpenStudy (anonymous):

that example you gave is just converting rev.s to rads.

OpenStudy (anonymous):

exactly....I will see if there's something like it in a different...sect.

OpenStudy (anonymous):

nm.

OpenStudy (anonymous):

Well, that's just a question at the end of the chap....just case something shows up like it on the exam. We didn't do a prob. like that so I better move on...we did do plenty of other problems to work on.

OpenStudy (anonymous):

here: (2pi radians) *1/2 m = circumference (one rev) so: find V in meters/min r*theta/min = V

OpenStudy (anonymous):

But, I don't know theta.

OpenStudy (anonymous):

you know V and r

OpenStudy (anonymous):

theta/min = V/r

OpenStudy (anonymous):

you're over thinking this... if a rolling wheel makes one complete rev (2pi radians), it has traveled forward a distance of 2pi*r two rev.s -> 4pi*r etc so whatever distance it travels forward = theta*r in a minute this wheel travels 1750m so theta*r =1750 theta = 3500 rad.s aka V/r

OpenStudy (anonymous):

ah....that's easier...

OpenStudy (anonymous):

I guess I don't understand how to find theta.

OpenStudy (anonymous):

think about a string wrapped around the rim of a tire. the string is as long as the circumference. roll the tire forward one complete rev. the string unrolls completely: the tire has traveled forward one circumference = 2pi*r

OpenStudy (anonymous):

it's 3500 rad.s every minute.

OpenStudy (anonymous):

k, good on this? is it clearer? or questions?

OpenStudy (anonymous):

I think am I ok...It's a lot to cram in my simple brain. lol. I am trying...:)

OpenStudy (anonymous):

what are you studying anyway?

OpenStudy (anonymous):

some physics+math class without calculus?

OpenStudy (anonymous):

or two different classes?

OpenStudy (anonymous):

trig and Microbiology